# Dipole-dipole scattering: summing large Pomeron loops in non-linear evolution with leading twist kernel

### v1

@@ I Introduction
where $z=\ln\left(r^{2}Q^{2}_{s}(Y)\right)$ [^2] We will discuss the definition of $z$ and of $\kappa$ in more detail below in Eq. (7) and in Eq. (9).. In our previous works we found that the sum of the large Pomeron loops leads to factor $\exp\left(-\frac{z^{2}}{\kappa}\right)$ in the scattering amplitudes of dipole-dipoles, dipole-nucleus and nucleus-nucleus. In this paper we are going to discuss the smooth function $C\left(z\right)$ in Eq. (1) for dipole-dipole scattering. However at the moment we can do this for the simplified BFKL kernel: the BFKL kernel in the leading twist.
@@ II Balitsky-Kovchegov non-linear equation for the leading twist BFKL kernel
$\phi_{\gamma}\left(\bm{r},\bm{R},\bm{b}\right)\,\,\,=\,\,\,\left(\frac{r^{2}\,R^{2}}{\left(\bm{b}+\frac{1}{2}(\bm{r}-\bm{R})\right)^{2}\,\left(\bm{b}-\frac{1}{2}(\bm{r}-\bm{R})\right)^{2}}\right)^{\gamma}=e^{\gamma\,\xi_{r,R}}~~\mbox{with}\,\,0\,<\,Re\,gamma\,<\,1$ (24)
@@ IV BK scattering amplitude in the BFKL Pomeron calculus
$A^{dip}_{A}\left(Y,r,R;\bm{b}\right)=\sum_{n=1}^{\infty}C_{n}(-1)^{n}\prod^{n}_{i=1}\int\frac{d^{2}r_{i}}{r^{2}_{i}}\,G_{I\!\!P}\left(Y\xi_{r}\right)\gamma^{BA}_{A}\left(r_{i},b\right)G_{I\!\!P}\left(Y_{0},\xi_{r^{\prime}_{1},R}\right)$ (39)
@@
We are looking for the solution in the form: $c_{n}\,\,=\,\,\phi\left(n\right)\exp\left(\frac{1}{2}n^{2}\right)$ where $\phi\left(n\right)$ is a smooth function of $n$ in comparison with $\exp\left(\frac{1}{2}n^{2}\right)$. With such $c_{n}$ the integral over $k$ in Eq. (44) can be taken using the method of steepest decent, since the integrant has to strong maxima at the limits of integrations: $k=1$ and k=n1k=n-1. Introducing a new variable $k=1+\delta k$ in the first maximum, one can see that the main contribution comes from small $\delta k\sim 1/n$ at large $n$. Rewriting $\frac{1}{2}\left((n-k)+k^{2}\right)$as $\frac{1}{2}(n-1)^{2}-(n-1)\delta k+2\delta k$ one can see that the equation takes the form at large $n$:
@@
$N^{dip}_{A}\left(Y,r,R;\bm{b}\right)\,\,=\,\,1\,\,\,-\,\,\exp\left(-\frac{1}{8}z^{2}_{A}+z_{A}\,-\,\ln\left(\frac{1}{2}z_{A}\right)\right)$ (51)
@@
$\Omega\left(z_{A}\right)\,=\,\,\frac{1}{8}z^{2}_{A}-z_{A}\,+\,\ln\left(\frac{1}{2}z_{A}\right)$ (52)
In Fig. 8 we compare this solution with the exact numerical solution of Eq. (42). One can see that we did not find a good approximation for large $z$. .
Figure 8: $\Omega\left(z\right)$ versus $z$. $\Omega_{exact}$ is the numerical solution of Eq. (26) with $\Omega_{0}=0.1$. $\Omega\left(z\right)\,=\,\,\frac{1}{8}z^{2}_{A}-z_{A}\,+\,\ln\left(\frac{1}{2}z_{A}\right)$ is our approximation for $\Omega$ that is coming from the large $n$ behaviour of the coefficient $C_{n}$ in the Pomeron expansion of Eq. (48). $\Omega\left(z\right)\,=\,\,\frac{1}{8}z^{2}_{A}+\frac{1}{2}z_{A}\,+\,\ln\left(\frac{1}{2}z_{A}\right)$ is our choice of $C_{n}$ of Eq. (53). $\Omega=z^{2}/8$ is the asymptotic solution of Ref.LETU .
@@
with arbitrary coefficients $\beta$ and $\alpha$ is also the solution. For $\alpha=-1/2$ and $\beta=1$ we obtain a good description of $\Omega$( see Fig. 8), making $C_{n}$ of Eq. (53) with $\alpha=-1/2$ and $\beta=1$ our preferable choice.
@@ VII Multiplicity distribution of produced gluons for dipole-dipole scattering
$\,=\,$ $\,\frac{1}{2\sqrt{2}}\,\int\limits^{\infty}_{-\ln N\left(z^{\prime}\right)}\!\!\!\!\!\!\!\!\!d\lambda\,\underbrace{e^{-\frac{\lambda^{2}}{4}}\frac{1}{e^{\lambda}\,N\left(z^{\prime}\right)\sqrt{2\sqrt{\frac{\xi}{e^{\lambda}}}}}\,\exp\Bigg(-2\sqrt{\frac{\xi}{e^{\lambda}}}\Bigg)}_{exp.}$ (81)
Fig. 11 shows the $\lambda$ dependence of $II\left(\lambda,\xi\right)$ . One can see that (i) this dependence has a clear maximum; (ii) the lower limit of integration does not influence on the value of the integral; and (iii) the approximate expression for $II$ ( see exp. in Eq. (VII)) describe the exact integrant quite well. Therefore, we can use the last term in Eq. (VII) and take the integral using the method of steepest descent with the following equation for $\lambda_{SP}$:
@@ IX Conclusions
In this paper we found the dipole densities (see Eq. (67)) for the leading twist BFKL kernel. These densities stem from the solution of nonlinear Balitsky-Kovchegov evolution equation for the scattering amplitude, and $t$-channel unitarity of Eq. (2) with the simple Eq. (38) for the dipole densities of a nucleus. Recall that Eq. (38) views the nucleus as a bag of dipoles. Having these densities we use Eq. (2) to calculate the contributions of the large Pomeron loops to the scattering amplitude of dipole - dipole interactions at high energies.
@@
We have to remind that $C_{n}$ of Eq. (67) have been found for large $n$ (see Eq. (47) and Appendix). As we have shown in the appendix we could rely on their value for $z\geq 5$.
We wish to point out that our amplitude give the same asymptotic behaviour at large $z$ as was calculated in Ref.LETU ; IAMU . This result not only confirms the general idea of Ref.IAMU that ’rare’ fluctuation contributes to the high energy behaviour of the scattering amplitude but it gives a possibility to calculate its value in the approach based on the BFKL Pomeron calculus.
@@ Acknowledgements
Appendix A Alternative calculation of BK amplitude and values of $C_{n}$
In this appendix we return to the calculation of BK amplitude and are going to use Eq. (LABEL:SABA10) but insert $C_{n}$ in the form of Eq. (53). In addition we take off. the factor $1/n$ as it has discussed in section IV. Finally the equation takes the form:
$N^{dip}_{A}\left(Y,r,R;\bm{b}\right)=$ (87)
$\oint_{C}\frac{dn}{2\,\pi\,i}\exp\left(\frac{1}{2}n^{2}-2\,\alpha\,n\right)\frac{\pi}{\sin\left(\pi\,n\right)}G^{n}_{I\!\!P}\left(Y,\xi_{r,R},b\right)=\frac{1}{\sqrt{2\,\pi}}\int\limits^{\infty}_{-\infty}d\lambda\,e^{-h\lambda^{2}}\,\oint_{C}\frac{dn}{2\,\pi\,i}\frac{\pi}{\sin\left(\pi\,n\right)}e^{n\left(\lambda\,+\,\frac{1}{2}\,\,(z_{A}-4\alpha)\right)}$
Moving contour $C$ to position $C^{\prime\prime}$ in Fig. 7 and tasking into account that at small $n$ we can replace $\pi/\sin\left(\pi\,n\right)$ by $1/n$ we obtain:
$N^{dip}_{A}\left(Y,r,R;\bm{b}\right)\,=\,\frac{1}{\sqrt{2\,\pi}}\int\limits^{\infty}_{-\infty}d\lambda\,e^{-h\lambda^{2}}\,\oint_{C^{\prime\prime}}\frac{dn}{2\,\pi\,i}\frac{1}{n}\exp\Bigg(n\underbrace{\left(\lambda\,+\,\frac{1}{2}\,\,(z_{A}-4\alpha)\right)}_{t}\Bigg)$
Integral over $n$ gives $\Theta\left(t\right)$ and Eq. (A) reduces to
$N^{dip}_{A}\left(Y,r,R;\bm{b}\right)\,=\,\frac{1}{\sqrt{2\,\pi}}\int\limits^{\infty}_{-h(z_{A}-4\alpha)}\!\!\!\!\!\!\!\!\!d\lambda\,e^{-h\lambda^{2}}\,=1\,-\,\pi\,\text{Erfc}\left(\frac{z_{A}-4\alpha}{2\sqrt{2}}\right)$ (88)
Refer to caption Refer to caption
Fig. 12-a Fig. 12-b
Figure 12: $\Omega\left(z^{\prime}\right)$ and the scattering amplitudes $N\left(z^{\prime}\right)$ versus $z_{A}$ for BK amplitude. $C_{n}=\exp\left(\frac{1}{2}n^{2}+3.04\,n\right)$ are taken from Eq. (88).
Fig. 12 shows that we can describe $\Omega$ much better if we choose $\alpha=-0.76$. Therefore, Eq. (88) can describe the scattering amplitude quite well for $z_{A}\geq 5$.
It has to be mentioned that $C_{n}=\exp\left(\frac{1}{2}n^{2}+3.04\,n\right)$ does not change any of our results but allow us to hope that the dipole-dipole scattering amplitude as well as the multiplicity distributions can be trusted from rather small $z\approx 5$.

### v2

@@ I Introduction
where $z=\ln\left(r^{2}Q^{2}_{s}(Y)\right)$ [^2] We will discuss the definition of $z$ and of $\kappa$ in more detail below in Eq. (7) and in Eq. (9).. In our previous works we found that the sum of the large Pomeron loops leads to factor $\exp\left(-\frac{z^{2}}{4\,\kappa}\right)$ in the scattering amplitudes of dipole-dipoles, dipole-nucleus and nucleus-nucleus. In this paper we are going to discuss the smooth function $C\left(z\right)$ in Eq. (1) for dipole-dipole scattering. However at the moment we can do this for the simplified BFKL kernel: the BFKL kernel in the leading twist.
@@ II Balitsky-Kovchegov non-linear equation for the leading twist BFKL kernel
$\phi_{\gamma}\left(\bm{r},\bm{R},\bm{b}\right)\,\,\,=\,\,\,\left(\frac{r^{2}\,R^{2}}{\left(\bm{b}+\frac{1}{2}(\bm{r}-\bm{R})\right)^{2}\,\left(\bm{b}-\frac{1}{2}(\bm{r}-\bm{R})\right)^{2}}\right)^{\gamma}=e^{\gamma\,\xi_{r,R}}~~\mbox{with}\,\,0\,<\,Re\,\gamma\,<\,1$ (24)
@@ IV BK scattering amplitude in the BFKL Pomeron calculus
$A^{dip}_{A}\left(Y,r,R;\bm{b}\right)=\sum_{n=1}^{\infty}C_{n}(-1)^{n}\prod^{n}_{i=1}\int\frac{d^{2}r_{i}}{r^{2}_{i}}\,G_{I\!\!P}\left(Y,\xi_{r}\right)\gamma^{BA}_{A}\left(r_{i},b\right)G_{I\!\!P}\left(Y_{0},\xi_{r^{\prime}_{1},R}\right)$ (39)
@@
We are looking for the solution in the form: $c_{n}\,\,=\,\,\phi\left(n\right)\exp\left(\frac{1}{2}n^{2}\right)$ where $\phi\left(n\right)$ is a smooth function of $n$ in comparison with $\exp\left(\frac{1}{2}n^{2}\right)$. With such $c_{n}$ the integral over $k$ in Eq. (44) can be taken using the method of steepest decent, since the integrant has to strong maxima at the limits of integrations: $k=1$ and k=n1k=n-1. Introducing a new variable $k=1+\delta k$ in the first maximum, one can see that the main contribution comes from small $\delta k\sim 1/n$ at large $n$. Rewriting $\frac{1}{2}\left((n-k)^{2}+k^{2}\right)$as $\frac{1}{2}(n-1)^{2}-n\delta k+2\delta k^{2}$ one can see that the equation takes the form at large $n$:
@@
$N^{dip}_{A}\left(Y,r,R;\bm{b}\right)\,\,=\,\,1\,\,\,-\,\,\exp\left(-\frac{1}{8}z^{2}_{A}+z_{A}\,-\,\ln\left(\frac{1}{2}z_{A}\right)-2\right)$ (51)
@@
$\Omega\left(z_{A}\right)\,=\,\,\frac{1}{8}z^{2}_{A}-z_{A}\,+\,\ln\left(\frac{1}{2}z_{A}\right)\,+\,2$ (52)
In Fig. 8 we compare this solution with the exact numerical solution of Eq. (42). One can see that we found a good approximation for $z\geq 4$.
Refer to caption Refer to caption
Fig. 8-a Fig. 8-b
Figure 8: $\Omega\left(z\right)$ versus $z$. $\Omega_{exact}$ is the numerical solution of Eq. (26) with $\Omega_{0}=0.1$. $\Omega\left(z\right)\,=\,\,\frac{1}{8}z^{2}_{A}-z_{A}\,+\,\ln\left(\frac{1}{2}z_{A}\right)$ is our approximation for $\Omega$ that is coming from the large $n$ behaviour of the coefficient $C_{n}$ in the Pomeron expansion of Eq. (53). $\Omega=z^{2}/8$ is the asymptotic solution of Ref.LETU .
@@
with arbitrary coefficients $\beta$ and $\alpha$ is also the solution. Therefore, $\alpha=\beta=1$ is our preferable choice. Since, these values lead to good descriptions at small values of $z_{A}$ (see Fig. 8-b) we can hope that our $C_{n}$ could describe the region of small $n$. It is instructive to note that $z_{A}+2$ term in Eq. (52) comes from the choice $\alpha=1$ while $\ln(z_{A}/2)$ term reflects $1/n$ behaviour of $C_{n}$.
@@ VII Multiplicity distribution of produced gluons for dipole-dipole scattering
$\,=\,$ $\,\frac{1}{2\sqrt{2}}\,\int\limits^{\infty}_{-\ln N\left(z^{\prime}\right)}\!\!\!\!\!\!\!\!\!d\lambda\,\underbrace{e^{-\frac{\lambda^{2}}{4}}\frac{1}{e^{\lambda}\,N\left(z^{\prime}\right)\sqrt{2\sqrt{\frac{\xi}{e^{\lambda}}}}}\,\exp\Bigg(-2\sqrt{\frac{\xi}{e^{\lambda}}}\Bigg)}_{approx.}$ (81)
Fig. 11 shows the $\lambda$ dependence of $II\left(\lambda,\xi\right)$ . One can see that (i) this dependence has a clear maximum; (ii) the lower limit of integration does not influence on the value of the integral; and (iii) the approximate expression for $II$ ( see approx. in Eq. (VII)) describe the exact integrant quite well. Therefore, we can use the last term in Eq. (VII) and take the integral using the method of steepest descent with the following equation for $\lambda_{SP}$:
@@ IX Conclusions
In this paper we found the dipole densities (see Eq. (67)) for the leading twist BFKL kernel. These densities stem from the solution of nonlinear Balitsky-Kovchegov evolution equation for the scattering amplitude, and $t$-channel unitarity of Eq. (2) with the simple expression of Eq. (38) for the dipole densities of a nucleus. Recall that Eq. (38) views the nucleus as a bag of dipoles. Having these densities we use Eq. (2) to calculate the contributions of the large Pomeron loops to the scattering amplitude of dipole - dipole interactions at high energies.
@@
We have to remind that $C_{n}$ of Eq. (67) have been found for large $n$ (see Eq. (47) and Fig. 8-a). As we have shown (see Fig. 8-b) we could rely on their value for $z\geq 4$.
We wish to point out that our amplitude gives the same asymptotic behaviour at large $z$ as was calculated in Ref.LETU ; IAMU . This result not only confirms the general idea of Ref.IAMU that ’rare’ fluctuation contributes to the high energy behaviour of the scattering amplitude but it gives a possibility to calculate its value in the approach based on the BFKL Pomeron calculus.